# Robust estimation of boundary using doubly stochastic scaling of Gaussian kernel

### v2

@@ Abstract.
This paper addresses the problem of detecting points on or near the boundary of a dataset sampled, potentially with noise, from a compact manifold with boundary. We extend recent advances in doubly stochastic scaling of the Gaussian heat kernel via Sinkhorn iterations to this setting. Our main contributions are: (a) deriving a characterization of the scaling factors for manifolds with boundary, (b) developing a boundary direction estimator, aimed at identifying boundary points, based on doubly stochastic kernel and local principal component analysis, and (c) demonstrating through simulations that the resulting estimates of the boundary points outperform the standard Gaussian kernel-based approach, particularly under noisy conditions.
Keywords. Manifold with boundary, doubly stochastic kernel, boundary detection.
@@ 1. Introduction
Given a set of points sampled, potentially with additive noise, from a compact manifold with boundary [19], we address the problem of identifying points on or near the boundary. The knowledge of such points is crucial for numerically solving partial differential equations with boundary conditions [26], unbiased kernel density estimation [23, 12, 15], constructing low-distortion local eigenmaps for regions near the boundary [13, 2, 3], and evaluating the performance of clustering algorithms [14] in preserving the structure within clusters.
@@
Estimating the boundary of a data manifold is a relatively underexplored area. Berry and Sauer [1] proposed an estimation strategy aimed at the general case of data sampled from compact Riemannian manifolds in arbitrary dimensions. They employed the standard kernel density estimator (KDE) along with a boundary direction estimator (BDE) to numerically solve for the distances of the points from the boundary. In this work, we demonstrate that their standard Gaussian kernel-based estimates are highly sensitive to noise in the data. Similarly, Calder, Park, and Slepčev [6] introduced a boundary estimator that replaces the Gaussian kernel in KDE and BDE with a binary cutoff kernel. However, their method is only designed for (flat) Euclidean domains, as confirmed by our experiments. Moreover, we demonstrate that their method is also sensitive to noise.
@@
A recent line of research [20, 18, 17, 8] investigated the concept of doubly stochastic scaling via Sinkhorn iterations [10] applied to the standard Gaussian heat kernel. These studies demonstrated that the doubly stochastic kernel leads to significant improvements in kernel density estimation on a closed manifold, and in approximating the eigenvectors of manifold Laplacian, especially in the presence of high-dimensional heteroskedastic and outlier-type noise. Additionally, Cheng and Landa [8] established convergence results for the bi-stochastically normalized Laplacian to a weighted manifold Laplacian and also demonstrated its resilience to outlier noise. This paper follows a similar approach by utilizing the doubly stochastic kernel to obtain robust boundary estimates for data points sampled from a manifold with noise. The contributions of our paper are as follows:
  1. (1)
@@
    Similar to the characterization by Landa and Cheng [17], of the scaling factors $\rho_{h}(\mathbf{x})$ that transform a Gaussian heat kernel $k_{h}(\mathbf{x},\mathbf{y})$ into a doubly stochastic kernel on a closed manifold, in this work, we derive a characterization of the scaling factors in the setting of a manifold with boundary. Our second-order characterization in Theorem 2.1 has the form:
    (1) $A_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}-B_{h}(x)\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})+\pi^{d}=\mathcal{O}(h^{2})$
@@
    where $d$ is the dimension of the manifold, $h$ is the Gaussian kernel bandwidth, and $A_{h}(\mathbf{x})$ and $B_{h}(\mathbf{x})$ depend on the distance $b_{\mathbf{x}}$ of $\mathbf{x}$ from the boundary, mean curvature induced by the boundary at $\mathbf{x}$ and the derivative of the sampling density $q(\mathbf{x})$ in the direction normal to the boundary at $\mathbf{x}$. Since the latter quantities are typically unknown for a given data manifold, for practical purposes, we obtain a relaxed first-order characterization in Corollary 2.4,
@@
    where $\zeta_{h}(\mathbf{x})$ depends only on the distance of $\mathbf{x}$ from the boundary i.e. on $b_{\mathbf{x}}$.
  2. (2)
    We present a BDE $\{\bm{\nu}_{i}\}_{1}^{n}$ by leveraging the doubly stochastic heat kernel $\mathbf{W}$ and local principal component analysis (PCA). Precisely,
@@
    where the columns of $\widetilde{\mathbf{U}}_{i}\in\mathbb{R}^{m\times d}$ represent the $d$ principal directions obtained by applying PCA in a neighborhood of the $i$-th point $\widetilde{\mathbf{x}}_{i}$. In Theorem 3.1, we prove the convergence of $\bm{\nu}_{i}$ to the normal direction to the boundary $\bm{\eta}_{\mathbf{x}_{i}}$ up to a global factor that depends only on the distance to the boundary. To keep the focus of the paper on doubly stochastic scaling, our convergence analysis assumes access to the projector onto the tangent space. In our experiments, this projector is approximated using local PCA on noisy data points. With this assumption, we show that the following holds with high probability:
    (4) $\bm{\nu}_{i}\xrightarrow[]{\ n\rightarrow\infty\ }h\beta_{h}(\mathbf{x})\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{2}\right).$
@@
The organization of the paper is as follows: in Section 1.1 we introduce the setup. In Section 2, we provide a characterization of the doubly stochastic scaling factors on manifolds with boundary. In Section 3, we present our boundary direction estimator and analyze its convergence. In Section 4, we showcase our experimental results demonstrating the effectiveness of our method, followed by conclusions in Section 5. Proofs of our results are provided in Section 6.
1.1. Setup and Preliminaries
@@
Suppose $(\mathbf{x}_{i})_{1}^{n}\subset\mathbb{R}^{m}$ are sampled independently and identically from a smooth and positive probability distribution $q(\mathbf{x})$ on $\mathcal{M}$. Let $\widetilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\bm{\varepsilon}(\mathbf{x}_{i})$ be the observed samples where $\{\bm{\varepsilon}(\mathbf{x}_{i})\}_{1}^{n}\subset\mathbb{R}^{m}$ are independent, perhaps non-identical, sub-Gaussian random variables with zero mean and a sub-Gaussian norm of $\left\|\bm{\varepsilon}(\mathbf{x}_{i})\right\|_{\psi_{2}}$. The noiseless setting is characterized by the constraint $\varepsilon(\mathbf{x}_{i})=0$ for all $i\in{1,\ldots,n}$. For $h>0$, define the Gaussian kernel $k_{h}$ and the affinity matrix $\mathbf{K}_{ij}$ by
@@
(6) $\mathbf{K}_{ij}$$\coloneqq\begin{cases}k_{h}(\widetilde{\mathbf{x}}_{i},\widetilde{\mathbf{x}}_{j})&i\neq j\\ 0&i=j.\end{cases}$
Given the kernel matrix $\mathbf{K}$, the doubly stochastic matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$ is defined as in [20, 18] by
(7) $\mathbf{W}_{ij}=\mathbf{d}_{i}\mathbf{K}_{ij}\mathbf{d}_{j}\text{ and }\sum_{j=1}^{n}\mathbf{W}_{ij}=1.$
where $\mathbf{d}\in\mathbb{R}^{n}$ are scaling factors. The existence of $\mathbf{d}$ follows from [18] and these are obtained efficiently using Sinkhorn iterations [10]. In the continuous setting, these correspond to a positive function $\rho_{h}$ on $\mathcal{M}$ such that
[…]
The existence of a smooth $\rho_{h}$ on $\mathcal{M}$ that satisfy the above constraint follows from [20, 16, 4]. In this work, we only need the following assumption:
@@
Under the sub-Gaussian noise setup, Landa and Cheng in [17] established a connection between the doubly stochastic kernel obtained from noisy observations and the Gaussian kernel obtained from clean samples. We are going to use their result and the underlying assumptions, as stated below, in the convergence analysis of our BDE in Section 3.
Theorem 1.1 (Theorem 2.1 in [17]).
@@
  • (A5)
    There exists $\gamma>0$ such that $m\geq n^{\gamma}$ (this facilitates the analysis of doubly stochastic scaling in the large sample size and high dimension regime).
@@
where $\mathcal{E}_{i,j}=\mathcal{O}^{(h)}_{m,n}(g(m,n))$, $g(m,n)=g_{0}(n)$ in the noiseless setting and $g(m,n)=\max\{g_{0}(n),g_{1}(m)\}$ otherwise. Here,
(11) $g_{0}(n)$ $\coloneqq\sqrt{\log n/n}\ \ \text{ and }$
(12) $g_{1}(m)$ $\coloneqq\max\{m^{-1/4},1/\sqrt{\log m}\}$
and as defined in [17], a random variable $X=\mathcal{O}^{(h)}_{m,n}(g(m,n))$ if there exist $t_{0}$, $n_{0}(h)$, $m_{0}(h)$, $C(h)>0$ such that for all $n>n_{0}(h)$, $m\geq m_{0}(h)$, $\mathrm{Pr}\{|X|\leq tC(h)g(m,n)\}\geq 1-n^{-t}$ for all $t\geq t_{0}$.
@@
The above result holds for continuous density $q$. However, we make the following stronger assumption throughout this work.
@@ 2. Characterizing doubly stochastic scaling factors on manifolds with boundary
In [17], the authors characterized the scaling factors $\rho_{h}(\mathbf{x})$ as defined in Eq. (8) in the setting of a closed manifold. Here, we obtain a characterization of $\rho_{h}(\mathbf{x})$ on manifolds with boundary. First, we need the following lemmas: (a) on the third-order approximation of the integral against Gaussian kernel on a manifold with boundary [9, 25], and (b) on the directional derivative of the projection operator onto the tangent space.
Lemma 2.1 (Theorem 4.7 in [25]).
@@
[…]
(14) $\qquad\frac{h^{2}}{2}\widetilde{m}_{2}\left(\widetilde{\omega}(\mathbf{x})f(\mathbf{x})+\Delta f(\mathbf{x})+\left(\frac{m^{(2)}_{h}(\mathbf{x})}{\widetilde{m}_{2}}-1\right)\partial_{\bm{\eta}_{\mathbf{x}}}^{2}f(\mathbf{x})\right)+\mathcal{O}(h^{3})$
where $H(\mathbf{x})$ is the mean curvature of the hypersurface parallel to $\partial\mathcal{M}$ intersecting $\mathbf{x}$ (which depends on the second fundamental form of $\partial\mathcal{M}\subset\mathcal{M}$), $\widetilde{\omega}(\mathbf{x})$ depends on the scalar curvature of $\mathcal{M}$ and
(15) $m_{\varepsilon}^{(0)}(\mathbf{x})$ $\coloneqq\frac{\pi^{d/2}}{2}\left(1+\mathrm{erf}\left(\frac{b_{\mathbf{x}}}{\varepsilon}\right)\right)$
(16) $m_{\varepsilon}^{(1)}(\mathbf{x})$ $\coloneqq-\frac{\pi^{(d-1)/2}}{2}e^{-b_{\mathbf{x}}^{2}/\varepsilon^{2}}$
(17) $m^{(2)}_{\varepsilon}(\mathbf{x})$ $\coloneqq\frac{b_{\mathbf{x}}}{\varepsilon}m^{(1)}_{\varepsilon}(\mathbf{x})+\frac{m^{(0)}_{\varepsilon}(\mathbf{x})}{2}$
[…]
In the following result, by applying Lemma 2.1 to Eq. 8, we derive the directional derivatives of the scaling factors $\rho_{h}(\mathbf{x})$ and a quadratic equation in $\rho_{h}^{2}(\mathbf{x})$ that characterizes the scaling factors on a manifold with boundary.
@@
Let $\mathcal{M}$ be a compact $d$-dimensional $\mathcal{C}^{3}$ Riemannian manifold with a $\mathcal{C}^{3}$ boundary and a normal collar. Suppose assumptions (A1) and (A2’) hold: $\rho_{h}\in\mathcal{C}^{3}(\mathcal{M})$ be a positive scaling function satisfying Eq. (8) and $q\in\mathcal{C}^{3}(\mathcal{M})$ be the positive sampling density. Then the following hold.
@@
(19) $h\partial_{\bm{\eta}_{\mathbf{x}}}\rho_{h}(\mathbf{x})=-\frac{2\rho_{h}(\mathbf{x})^{3}(m^{(1)}_{h}(\mathbf{x})q(\mathbf{x})+hm^{(2)}_{h}(\mathbf{x})\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x}))}{\pi^{d/2}+2m^{(2)}_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})}+\mathcal{O}(h^{2})$
@@
(II) $\rho_{h}(\mathbf{x})$ satisfies the following equation,
(21) $A_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}-B_{h}(x)\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})+\pi^{d}=\mathcal{O}(h^{2})$
where $A_{h}$ and $B_{h}$ are defined as follows.
(22) $A_{h}(\mathbf{x})=2\left\{m^{(1)}_{h}(\mathbf{x})^{2}-m^{(2)}_{h}(\mathbf{x})\left(m^{(0)}_{h}(\mathbf{x})+\frac{h(d-1)m^{(1)}_{h}(\mathbf{x})H(\mathbf{x})}{2}\right)\right\}\text{ and}$
(23) $B_{h}(\mathbf{x})=\pi^{d/2}m^{(1)}_{h}(\mathbf{x})\left(h\omega(\mathbf{x})-\frac{2b_{\mathbf{x}}}{h}\right),$
where $\omega(\mathbf{x})\coloneqq\frac{\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})}{q(\mathbf{x})}+\frac{(d-1)}{2}H(\mathbf{x})$ depends on the sampling density and the mean curvature $H(\mathbf{x})$ induced by the boundary at $\mathbf{x}$. Since $\mathcal{M}$ is compact, the constant in $\mathcal{O}(h^{2})$ can be uniformly bounded over the manifold.
The proof of the above theorem, when specialized to closed and compact manifolds, results in the following corollary, which is reminiscent of the characterization of scaling factors presented in [17].

### v3

# Robust boundary detection and density estimation using doubly stochastic scaling of the Gaussian kernel

@@ Abstract.
This paper addresses the problem of detecting boundary points and estimating the sampling density of a dataset derived from a compact manifold with boundary, potentially in the presence of noise. We extend recent advances in doubly stochastic scaling of the Gaussian heat kernel via Sinkhorn iterations to this setting. Our main contributions are: (a) deriving a characterization of the scaling factors for manifolds with boundary, (b) developing a boundary direction estimator aimed at identifying boundary points followed by a boundary-corrected kernel density estimator based on doubly stochastic kernel and local principal component analysis, and (c) demonstrating through simulations that the resulting estimates of the boundary points and the sampling density outperform the standard Gaussian kernel-based approach, particularly under noisy conditions.
Keywords. Manifold with boundary, doubly stochastic kernel, boundary detection, kernel density estimation.
@@ 1. Introduction
We address the problem of identifying points on or near the boundary of a compact manifold [lee2018introduction] using a potentially noisy set of sampled points. The knowledge of such points is crucial for numerically solving partial differential equations with boundary conditions [VAUGHN2024101593], unbiased kernel density estimation [schuster1985incorporating, jones1996simple, karunamuni2005boundary], constructing low-distortion local eigenmaps for regions near the boundary [jones2007universal, ldle1v2, ratsv2], and evaluating the performance of clustering algorithms [jung2018network] in preserving the structure within clusters.
@@
Estimating the boundary of a data manifold remains a relatively underexplored problem. Berry and Sauer [berry2017density] proposed an estimation strategy aimed at the general case of data sampled from compact Riemannian manifolds in arbitrary dimensions. They employed the standard Gaussian kernel density estimator (KDE) along with a boundary direction estimator (BDE) to numerically solve for the distances of the points from the boundary. Similarly, Calder, Park, and Slepčev [calder2022boundary] introduced a boundary estimator that replaces the Gaussian kernel in the KDE and BDE with a binary cutoff kernel. In this work, we show that boundary estimates based on both the standard Gaussian kernel and the binary cutoff kernel are highly sensitive to noise. Moreover, the latter method is also sensitive to the curvature of the underlying manifold, as demonstrated in our experiments.
@@
A recent line of research [marshall2019manifold, landa2021doubly, landa2023robust, cheng2024bi] investigated the concept of doubly stochastic scaling via Sinkhorn iterations [cuturi2013sinkhorn] applied to the standard Gaussian heat kernel. These studies demonstrated that the doubly stochastic kernel leads to significant improvements in kernel density estimation on a closed manifold, and in approximating the eigenvectors of the manifold Laplacian, especially in the presence of high-dimensional heteroskedastic and outlier-type noise. Additionally, Cheng and Landa [cheng2024bi] established convergence results for the bi-stochastically normalized Laplacian to a weighted manifold Laplacian and also demonstrated its resilience to outlier noise. In this paper, we follow a similar approach by utilizing the doubly stochastic kernel to obtain robust boundary and density estimates for data points sampled from a manifold with noise. The contributions of our paper are as follows.
@@
First, building on the characterization by Landa and Cheng [landa2023robust] of the scaling factors $\rho_{h}(\mathbf{x})$ that transform a Gaussian heat kernel $k_{h}(\mathbf{x},\mathbf{y})$ with bandwidth $h$ into a doubly stochastic kernel on a closed manifold, we derive an analogous characterization of the scaling factors in the setting of a manifold with boundary. Our second-order characterization, presented in Theorem 3.1, takes the form:
(2) $A_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}-B_{h}(\mathbf{x}))\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})+\pi^{d}=\mathcal{O}(h^{2})$
@@
where $d$ is the dimension of the manifold. The coefficients $A_{h}(\mathbf{x})$ and $B_{h}(\mathbf{x})$ depend on the distance $b_{\mathbf{x}}$ of $\mathbf{x}$ from the boundary, the mean curvature induced by the boundary at $\mathbf{x}$, and the derivative of the sampling density $q(\mathbf{x})$ in the direction normal to the boundary at $\mathbf{x}$. Since these quantities are typically unknown for a given data manifold, for practical purposes, we obtain a relaxed first-order characterization in Corollary 3.4,
@@
where $\zeta_{h}(\mathbf{x})$ depends only on the distance of $\mathbf{x}$ from the boundary, i.e. on $b_{\mathbf{x}}$.
Second, given a noisy point cloud $\{\widetilde{\mathbf{x}}_{i}\}_{1}^{n}$, we introduce a BDE $\{\bm{\nu}_{i}\}_{1}^{n}$ based on the doubly stochastic heat kernel $\mathbf{W}$ [marshall2019manifold, landa2021doubly, landa2023robust, cheng2024bi] and local principal component analysis (PCA) [cleveland1988locally, arias2011spectral, singer2012vector, tyagi2013tangent, cheng2013local, kaslovsky2014non, yu2015useful, mohammed2017manifold, aamari2019nonasymptotic, gilbert2025pca]. Specifically,
@@
where the columns of $\widetilde{\mathbf{U}}_{i}\in\mathbb{R}^{m\times d}$ represent the $d$ principal directions obtained by applying PCA in a neighborhood of the $i$-th noisy data point $\widetilde{\mathbf{x}}_{i}$. In Theorem 4.1, we establish the convergence of $\bm{\nu}_{i}$ to the normal direction to the boundary $\bm{\eta}_{\mathbf{x}_{i}}$ at $\mathbf{x}_{i}$, the clean counterpart of $\widetilde{\mathbf{x}}_{i}$, up to a global scaling factor that depends only on the distance to the boundary. To keep the focus of the paper on doubly stochastic scaling, our convergence analysis assumes access to the projector onto the tangent space. In our experiments, this projector is approximated via local PCA on noisy data points. Under this assumption, we show that the following holds with high probability:
(5) $\bm{\nu}_{i}\xrightarrow[]{\ n\rightarrow\infty\ }h\beta_{h}(\mathbf{x}_{i})\bm{\eta}_{\mathbf{x}_{i}}+\mathcal{O}\left(h^{2}\right).$
@@
Finally, we propose a two-parameter family of kernel density estimators for manifolds with boundary that corrects the boundary-induced bias inherent in standard KDEs [marron1994transformations, cowling1996pseudodata, karunamuni2005boundary] by explicitly incorporating distances to the boundary. This approach extends the robust KDE framework of [landa2023robust] from closed manifolds to manifolds with boundary while remaining effective under heteroskedastic noise.
The organization of the paper is as follows: in Section 2 we introduce the setup. In Section 3, we provide a characterization of the doubly stochastic scaling factors on manifolds with boundary. In Section 4, we present our BDE and KDE, and analyze their convergence. In Section 5, we showcase our experimental results demonstrating the effectiveness of our method, followed by conclusions in Section 6. Proofs of our results are provided in Appendix A.
2. Setup and Preliminaries
@@
Suppose $(\mathbf{x}_{i})_{1}^{n}\subset\mathbb{R}^{m}$ are sampled independently and identically from a smooth and positive probability distribution $q(\mathbf{x})$ on $\mathcal{M}$. Let $\widetilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\bm{\varepsilon}(\mathbf{x}_{i})$ be the observed noisy samples where $\{\bm{\varepsilon}(\mathbf{x}_{i})\}_{1}^{n}\subset\mathbb{R}^{m}$ are independent, perhaps non-identical, sub-Gaussian random variables with zero mean and a sub-Gaussian norm of $\left\|\bm{\varepsilon}(\mathbf{x}_{i})\right\|_{\psi_{2}}$. The noiseless setting is characterized by the constraint $\varepsilon(\mathbf{x}_{i})=0$ for all $i\in{1,\ldots,n}$. For $h>0$, define the Gaussian kernel $k_{h}$ and the affinity matrix $\mathbf{K}_{ij}$ by
@@
(7) $\mathbf{K}_{ij}$ $\coloneqq\begin{cases}k_{h}(\widetilde{\mathbf{x}}_{i},\widetilde{\mathbf{x}}_{j}),&i\neq j\\ 0,&i=j.\end{cases}$
Given the kernel matrix $\mathbf{K}$, the doubly stochastic matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$ is defined as in [marshall2019manifold, landa2021doubly] by
(8) $\mathbf{W}_{ij}=\mathbf{d}_{i}\mathbf{K}_{ij}\mathbf{d}_{j}\text{ and }\sum_{j=1}^{n}\mathbf{W}_{ij}=1,$
where $\mathbf{d}\in\mathbb{R}^{n}$ are scaling factors. The existence of $\mathbf{d}$ follows from [landa2021doubly] and these are obtained efficiently using Sinkhorn iterations [cuturi2013sinkhorn]. In the continuous setting, these correspond to a positive function $\rho_{h}$ on $\mathcal{M}$ such that
(9) $\frac{1}{\pi^{d/2}h^{d}}\int_{\mathcal{M}}\rho_{h}(\mathbf{x})k_{h}(\mathbf{x},\mathbf{y})\rho_{h}(\mathbf{y})q(\mathbf{y})dV(\mathbf{y})=1.$
The existence of a smooth $\rho_{h}$ on $\mathcal{M}$ that satisfies the above constraint follows from [marshall2019manifold, knopp1968note, borwein1994entropy]. In this work, we only need the following assumption:
@@
Under the sub-Gaussian noise setup, Landa and Cheng in [landa2023robust] established a connection between the doubly stochastic kernel obtained from noisy observations and the Gaussian kernel obtained from clean samples. We use their result and the underlying assumptions, as stated below, in the convergence analysis of our BDE in Section 4.
Lemma 2.1 (Theorem 2.1 in [landa2023robust]).
@@
  4. (A5)
    There exists $\gamma>0$ such that $m\geq n^{\gamma}$.
@@
where $\mathcal{E}_{i,j}=\mathcal{O}^{(h)}_{m,n}(g(m,n))$. In the noiseless case, $g(m,n)=g_{0}(n)$, whereas in the presence of noise, $g(m,n)=\max\{g_{0}(n),g_{1}(m)\}$. Here,
(12) $g_{0}(n)$ $\coloneqq\sqrt{\log n/n},$
(13) $g_{1}(m)$ $\coloneqq\max\{m^{-1/4},1/\sqrt{\log m}\},$
and as defined in [landa2023robust], a random variable $X=\mathcal{O}^{(h)}_{m,n}(g(m,n))$ if there exist $t_{0}$, $n_{0}(h)$, $m_{0}(h)$, $C(h)>0$ such that for all $n>n_{0}(h)$, $m\geq m_{0}(h)$,
$\mathrm{Pr}\{|X|\leq tC(h)g(m,n)\}\geq 1-n^{-t}\text{ for all }t\geq t_{0}.$
Moreover,
  1. (i)
    if $Y=f^{(h)}(X)$ where $f^{(h)}\in\mathcal{C}^{1}(\mathbb{R})$ for every $h>0$ then
    (14) $Y=f^{(h)}(0)+\mathcal{O}^{(h)}_{m,n}(g(m,n)).$
  2. (ii)
    if $X_{i}=\mathcal{O}^{(h)}_{m,n}(g(m,n))$ for all $i\in[1,n]$ then
    (15) $\max_{1}^{n}|X_{i}|=\mathcal{O}^{(h)}_{m,n}(g(m,n)).$
@@
Assumption (A4) bounds the noise in terms of the ambient dimension while the assumption (A5) controls the growth of the ambient dimension with the sample size. Overall, the two assumptions facilitate the analysis of doubly stochastic scaling in the large sample size and high dimension regime.
Remark 2.2.
The above result holds for continuous density $q$ on $\mathcal{M}$. However, we make the following stronger assumption throughout this work.
@@ 3. Characterizing doubly stochastic scaling factors on manifolds with boundary
In [landa2023robust], the authors characterized the scaling factors $\rho_{h}(\mathbf{x})$ as defined in Eq. (9) in the setting of a closed manifold. Here, we obtain a characterization of $\rho_{h}(\mathbf{x})$ on manifolds with boundary. Specifically, we obtain a second order characterization of $\rho_{h}(\mathbf{x})$ that depends on the sampling density $q(\mathbf{x})$, the mean curvature $H(\mathbf{x})$ induced by the boundary and the distance $b_{\mathbf{x}}$ to the boundary. Since, for a given dataset, the curvature $H(\mathbf{x})$ is not known a priori, we further obtain a first order characterization of $\rho_{h}(\mathbf{x})$ that only depends on $q(\mathbf{x})$ and $b_{\mathbf{x}}$ and later utilize this characterization to obtain robust estimate of points on the boundary.
First, we need the following lemmas: (a) on the third-order approximation of the integral against Gaussian kernel on a manifold with boundary [coifman2006diffusion, vaughn2019diffusion] as stated below, and (b) on the directional derivative of the projection operator onto the tangent space (see Lemma A.1 in the Appendix).
Lemma 3.1 (Theorem 4.7 in [vaughn2019diffusion]).
@@
(16) $\frac{1}{h^{d}}\int_{\mathcal{M}}k_{h}(\mathbf{x},\mathbf{y})f(\mathbf{y})dV(\mathbf{y})$
(17) $\hskip 21.33955pt=m^{(0)}_{h}(\mathbf{x})f(\mathbf{x})+hm^{(1)}_{h}(\mathbf{x})\left(\partial_{\bm{\eta}_{\mathbf{x}}}f(\mathbf{x})+\frac{d-1}{2}H(\mathbf{x})f(\mathbf{x})\right)+\frac{h^{2}}{2}\left\{\frac{\pi^{d/2}}{2}\widetilde{\omega}(\mathbf{x})f(\mathbf{x})+\right$.
(18) $\hskip 42.67912pt\left.\frac{\pi^{d/2}}{2}(\Delta f(\mathbf{x})-\partial_{\bm{\eta}_{\mathbf{x}}}^{2}f(\mathbf{x}))+m^{(2)}_{h}(\mathbf{x})(\partial_{\bm{\eta}_{\mathbf{x}}}^{2}f(\mathbf{x})-(d-1)H(\mathbf{x})\partial_{\eta_{\mathbf{x}}}f(\mathbf{x}))\right\}+\mathcal{O}(h^{3})$
where $H(\mathbf{x})$ is the mean curvature of the hypersurface parallel to $\partial\mathcal{M}$ intersecting $\mathbf{x}$ (which depends on the second fundamental form of $\partial\mathcal{M}\subset\mathcal{M}$), $\widetilde{\omega}(\mathbf{x})$ depends on the scalar curvature of $\mathcal{M}$ and the boundary, and
(19) $m_{h}^{(0)}(\mathbf{x})$ $\coloneqq\frac{\pi^{d/2}}{2}\left(1+\mathrm{erf}\left(\frac{b_{\mathbf{x}}}{h}\right)\right)$
(20) $m_{h}^{(1)}(\mathbf{x})$ $\coloneqq-\frac{\pi^{(d-1)/2}}{2}e^{-b_{\mathbf{x}}^{2}/h^{2}}$
(21) $m^{(2)}_{h}(\mathbf{x})$ $\coloneqq\frac{b_{\mathbf{x}}}{h}m^{(1)}_{h}(\mathbf{x})+\frac{m^{(0)}_{h}(\mathbf{x})}{2}.$
In the following result, by applying Lemma 3.1 and Lemma A.1 to Eq. 9, we derive the directional derivatives of the scaling factors $\rho_{h}(\mathbf{x})$ and a quadratic equation in $\rho_{h}^{2}(\mathbf{x})$ that characterizes the scaling factors on a manifold with boundary. These derivatives are utilized in Section 4 to establish the convergence of our BDE to the normal direction to the boundary, up to a scaling factor that depends on the distances to the boundary.
@@
Let $\mathcal{M}$ be a compact $d$-dimensional $\mathcal{C}^{3}$ Riemannian manifold with a $\mathcal{C}^{3}$ boundary and a normal collar. Suppose assumptions (A1) and (A2’) hold: $\rho_{h}\in\mathcal{C}^{3}(\mathcal{M})$ be a positive scaling function satisfying Eq. (9) and $q\in\mathcal{C}^{3}(\mathcal{M})$ be the positive sampling density. Then the following holds.
@@
(22) $h\partial_{\bm{\eta}_{\mathbf{x}}}\rho_{h}(\mathbf{x})=-\frac{2\rho_{h}(\mathbf{x})^{3}q(\mathbf{x})\left\{m^{(1)}_{h}(\mathbf{x})+hm^{(2)}_{h}(\mathbf{x})\kappa_{-}(\mathbf{x})\right\}}{\pi^{d/2}+2m^{(2)}_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})}+\mathcal{O}(h^{2}).$
@@
(II) $\rho_{h}(\mathbf{x})$ satisfies the equation,
(24) $A_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}-B_{h}(\mathbf{x})\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})+\pi^{d}=\mathcal{O}(h^{2})$
where $A_{h}$ and $B_{h}$ are defined as follows:
(25) $A_{h}(\mathbf{x})$ $=2\left\{m^{(1)}_{h}(\mathbf{x})^{2}-m^{(2)}_{h}(\mathbf{x})\left(m^{(0)}_{h}(\mathbf{x})+h(d-1)m^{(1)}_{h}(\mathbf{x})H(\mathbf{x})\right)\right\}\text{ and}$
(26) $B_{h}(\mathbf{x})$ $=\pi^{d/2}m^{(1)}_{h}(\mathbf{x})\left(h\kappa_{+}(\mathbf{x})-\frac{2b_{\mathbf{x}}}{h}\right),$
where $\kappa_{\pm}(\mathbf{x})\coloneqq\frac{\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})}{q(\mathbf{x})}\pm\frac{(d-1)}{2}H(\mathbf{x})$ depends on the sampling density and the mean curvature $H(\mathbf{x})$ induced by the boundary at $\mathbf{x}$. Since $\mathcal{M}$ is compact, the constant in $\mathcal{O}(h^{2})$ can be uniformly bounded over the manifold.
All the corollaries presented below are extensions of theorem 3.1. By specializing the proof of this theorem to closed and compact manifolds, we obtain the following corollary, which is reminiscent of the characterization of scaling factors presented in [landa2023robust].